# Internal diffraction dynamics of trilobite molecules

### v1

@@ III.3 Prototypical experimental setup
With this in mind, let us propose an alternate experimental scheme that utilizes a time-resolved pump probe methodology to monitor dynamical effects in real time. The basic idea is to utilize lower-lying low-$l$ ULRM to create the initial state and as a probe to monitor the system. Figure 6(a) illustrates a prototypical example of the proposed scheme. The initial state preparation is accomplished by a two photon transition to the Gaussian-like vibrational ground-state of the |$\ket{n=50,l=2}$ ULRM state. The vibrational ground-states for different principle quantum numbers feature a convenient range of mean positions and widths, providing the necessary flexibility needed to observe the dynamical effects we seek. They can also be conveniently used as a probe-state localized near a range of desired $R$-values. As an example, the ULRM is excited from the |$\ket{n=50,l=2}$ state to the trilobite electronic state using a microwave pump pulse, utilising $f$-state admixture of the high-$l$ molecule [30]. The microwave excitation is fast enough to consider it projective, preserving the vibrational wavepacket as the initial state that porpagates in the trilobite potential. A time-delayed probe-pulse then enables us to project the time evolving wavepacket onto a |$\ket{n=40,l=2}$ probe-state and obtain the time resolved correlation function. Figure 6(b)shows the predicted result of this proposed experiment. The train of evenly spaced diffraction peaks produced by scattering of the wavepacket from the trilobite potential is observed as a series of peaks in the time resolved correlation function. Rather conveniently, this framework allows for the initial state preparation, time propagation, and observation of quantum interference and diffraction effects all encompassed within a single molecular system.
We have ignored the effects of $p$-wave shape resonances and spin-structure in our analysis. Our present work is a proof of methodology meant to bring attention and interest to the study of wavepacket dynamics of trilobite molecules. However, one can mitigate the complexity introduced by spin-structure to some extent by careful preparation and state selection of the hyperfine state of the ground-state atom. Furthermore, while the $p$-wave shape resonances will substantially alter the long time-scale dynamics, the initial time dynamics should still qualitatively exhibit the diffractive effects discussed. Of course, it is importing to note that these compilcations can be eliminated by selecting a different ground-state perturbing atom, such as Sr which has no low energy, p-wave electron-atom resonances and lacks any hyperfine structure.
@@
Figure 6: Pictorial scheme of the suggested pump-probe experiment. (a) The vibrational ground state of the |$\ket{n=50,l=2}$ ULRM is pumped to the $n$=55 trilobite potential and subsequently propagates therein. Following a time lack the wave packet is projected onto the desired |$\ket{n=40,l=2}$ probe state. (b) Observable time resolved correlation function calculated for the given setup.

### v2

@@ III.4 Prototypical experimental setup
With this in mind, let us propose an alternate experimental scheme that utilizes a time-resolved pump probe methodology to monitor dynamical effects in real time. The basic idea is to utilize lower-lying low-$l$ ULRM to create the initial state and as a probe to monitor the system. Figure 6(a) illustrates a prototypical example of the proposed scheme. The initial state preparation is accomplished by a two photon transition to the Gaussian-like vibrational ground-state of the |$\ket{n=45,l=2}$ ULRM state. The vibrational ground-states for different principle quantum numbers feature a convenient range of mean positions and widths, providing the necessary flexibility needed to observe the dynamical effects we seek. They can also be conveniently used as a probe-state localized near a range of desired $R$-values. Similar pump-probe schemes utilising low-$l$ Rydberg states were also utilized in the study of ion-atom scattering recently [69, 70]. As an example, the ULRM is excited from the |$\ket{n=45,l=2}$ state to the $n$=55 trilobite electronic state using a microwave pump pulse, utilising $f$-state admixture of the high-$l$ molecule [30]. The microwave excitation is fast enough to consider it projective, preserving the vibrational wavepacket as the initial state that porpagates in the trilobite potential. A time-delayed probe-pulse then enables us to project the time evolving wavepacket onto a |$\ket{n=35,l=2}$ probe-state and obtain the time resolved correlation function. Figure 6(b)shows the predicted result of this proposed experiment for three different $n$-values. In each case the |$\ket{n-10,l=2}$ pump-state is excited to the |$\ket{n,trilobite}$, with |$\ket{n-20,l=2}$ being the probe-state. The state selection is done to facilitate comparable propagation times and correlation functions, as seen in Fig. 6(b). The train of evenly spaced diffraction peaks produced by scattering of the wavepacket from the trilobite potential is observed as a series of peaks in the time resolved correlation function. Note that the amplitudes of the correlation function, which carries the phase-information of the diffraction pattern, decreases with increasing $n$. Although not a systematic study of the $n$-dependence of the diffraction effect, the decrease in the amplitude of correlation function is consistent with our analysis that higher $n$-values exhibit reduced back-scattering and consequently display diminished diffraction effect. Rather conveniently, the temporal resolution achieved here is only limited by the time delay between the pump and probe pulses and their respective widths, both being finely tunable in the nanosecond regime. In comparison, direct spatio-temporal imaging as achieved in state-of-the-art ion microscopes [71, 72] limits the resolution to the order of $\sim 100$ nm and $\sim 10$ ns, rendering the desired phenomenon challenging to observe. Moreover, recent advances in the construction of long-range molecular potentials with tunable electronic character [73], allow the possibility of engineering initial and probe states that can be optically excited to a trilobite state of choice. Note that although the theoretical description of a pump-probe scheme is rather straightforward (and abundant in conventional molecular physics), a corresponding experimental realization of the scheme in ultracold systems is not without due challenges [69, 70]. But, within the scope of this work the pump-probe framework allows for the initial state preparation, time propagation, and observation of quantum interference and diffraction effects all encompassed within a single molecular system.
@@
Figure 6: Pictorial scheme of the suggested pump-probe experiment. (a) The vibrational ground state of the |$\ket{n=45,l=2}$ ULRM is pumped to the $n$=55 trilobite potential and subsequently propagates therein. Following a time lag the wave packet is projected onto the desired |$\ket{n=35,l=2}$ probe state. (b) Observable time resolved correlation function calculated for multiple $n$-values using the given setup. For each |$\ket{n,trilobite}$ PEC, the pump and probe states are selected to be |$\ket{n-10,l=2}$ and |$\ket{n-20,l=2}$ respectively, where $n\in\{55,56,57\}$.